# DiRW: Path-Aware Digraph Learning for Heterophily

### v1

@@ Abstract.
Recently, graph neural network (GNN) has emerged as a powerful representation learning tool for graph-structured data. However, most approaches are tailored for undirected graphs, neglecting the abundant information embedded in the edges of directed graphs (digraphs). In fact, digraphs are widely applied in the real world (e.g., social networks and recommendations) and are also confirmed to offer a new perspective for addressing topological heterophily challenges (i.e., connected nodes have complex patterns of feature distribution or labels). Despite recent significant advancements in DiGNNs, existing spatial- and spectral-based methods have inherent limitations due to the complex learning mechanisms and reliance on high-quality topology, leading to low efficiency and unstable performance. To address these issues, we propose Directed Random Walk (DiRW), which can be viewed as a plug-and-play strategy or an innovative neural architecture that provides a guidance or new learning paradigm for most spatial-based methods or digraphs. Specifically, DiRW incorporates a direction-aware path sampler optimized from the perspectives of walk probability, length, and number in a weight-free manner by considering node profiles and topological structure. Building upon this, DiRW utilizes a node-wise learnable path aggregator for generalized messages obtained by our proposed adaptive walkers to represent the current node. Extensive experiments on 9 datasets demonstrate that DiRW: (1) enhances most spatial-based methods as a plug-and-play strategy; (2) achieves SOTA performance as a new digraph learning paradigm.
@@ 1. Introduction
Graph neural networks (GNNs) have been widely used across node- (Wu et al., 2019; Wang and Leskovec, 2020; Li et al., 2024b), link- (Zhang and Chen, 2018; Tan et al., 2023; Li et al., 2024a), and graph-level tasks (Liang et al., 2023; Luo et al., 2023; Thost and Chen, 2021) and achieve satisfactory performance. Therefore, this graph-based deep learning technique holds great potential for applications, such as recommendation (Su et al., 2024; Yang et al., 2023; Cai et al., 2023), financial analysis (Balmaseda et al., 2023; Hyun et al., 2023; Qiu, 2023), and healthcare (Pfeifer et al., 2022; Ahmed et al., 2022). Despite their effectiveness, existing methods often overlook edge direction in natural graphs, leading to inevitable information loss and limited performance upper bound.
@@
Compared to undirected representation, digraphs are crucial for modeling the complex topologies present in the real world (e.g., web flow monitoring and bioinformatics), capturing more intricate node relationships. Additionally, the recently proposed A2DUG (Maekawa et al., 2023), Dir-GNN (Rossi et al., 2023) and ADPA (Sun et al., 2023) reveal a key insight: Edge direction offers a new perspective for effectively addressing the topological heterophily challenges that plague graph learning. Despite the growing attention towards DiGNNs, this field is still in its infancy and faces the following inherent limitations. (1) Spatial-based methods often stack multiple convolution layers and employ two independent sets of learnable parameters to capture directed information concealed in the out-edges and in-edges, resulting in over-smoothing concerns (Yan et al., 2022) and high computational costs (He et al., 2022; Kollias et al., 2022; Zhou et al., 2022). (2) Spectral-based methods heavily rely on high-quality directed topology (Zhang et al., 2021a; Tong et al., 2020a), and without this, extreme eigenvalues inevitably lead to sub-optimal performance. Meanwhile, strict theoretical assumptions affect practicality, which hinders their effective deployment in complex real-world scenarios. Therefore, it is urgent to develop a more efficient paradigm for digraph learning.
@@
To enhance the usability of DiGNNs, this paper focuses on spatial-based methods and proposes a novel (directed) path-based learning mechanism (Di)PathGNNs. As we all know, the entanglement of homophily and heterophily, where connected nodes exhibit intricate feature distributions and labels, has recently posed a significant challenge (Ma et al., 2021; Luan et al., 2022; Zheng et al., 2022; Platonov et al., 2023). Researchers strive to achieve robust learning within this complex topology. In this context, compared to the traditional message passing (i.e., neighbor expansion with strict spatial symmetry that disregards edge direction), we highlight the following advantages of DiRW to further clarify the motivation of our study: (1) Edge Direction and Node Order. The core of DiPathGNNs lies in performing well-designed random walks for each node, treating the paths as node-wise sequences. Advantage: DiRW fully considers edge direction and preserves the order of nodes within the walking paths, which is crucial for capturing structural insights. (2) Adaptive Expansion of Node Receptive Fields. DiPathGNNs adaptively extend hop-based neighbors to path-based neighbors by considering the characteristics of random walk, incorporating more homophilous signals. Advantage: DiRW ensures message aggregation among nodes with the same label, equivalent to data augmentation and highlighting label-specific positive signals for prediction. (3) Path-based Message Aggregation. Based on the above homophily-aware paths, DiPathGNNs aggregate these messages for predicting each node through a learnable mechanism. Advantage: DiRW facilitates the dense aggregation of direction-aware homophilous information, thus enhancing node prediction.
@@
Despite recent advancements in PathGNNs (Jin et al., 2022; Xie et al., 2023; Sun et al., 2022), their walk strategies are tailored for undirected graphs, lacking generalizability. Furthermore, the complex relationships inherent in directed topology pose unique challenges to the naive random walks, necessitating further investigations to develop fine-grained walk rules. To further elucidate, we present empirical analyses in Fig. 1(a)-2(b) that demonstrate these limitations and offer key insights.
@@
(1) L1: Neglect of edge directions. Existing PathGNNs often fail to account for the edge direction. Utilizing directed simple random walks (DiSRW) on digraphs inevitably leads to interruptions when encountering nodes with no outgoing edges. To investigate and visualize this limitation, we employ DiSRW and initiate a walk from each node on four digraphs. As the walk length increases, we record the proportion of complete paths out of the total sequences, as depicted in Fig. 1(a). To mitigate the impact of cycles within the digraph, we also devise an DiSRW that excludes cycles and conduct the same experiment, as shown in Fig. 1(b).
@@
Key Insight 1: Walking strictly along the edge directions leads to severe walk interruption problems. Due to the non-strongly connected nature of digraphs, most walks exhibit a sharp decline in the proportion of complete paths after only five steps. This indicates that the majority of walk sequences fail to gather extensive information beyond the immediate neighborhood of the originating node. Furthermore, when the influence of cycles is removed, the proportion of uninterrupted sequences declines more significantly.
@@
(2) L2: Coarse-grained walk strategies. Most PathGNNs treat walk number and length as hyperparameters, uniformly applying them across all nodes. This one-size-fits-all approach overlooks the distinct contexts of nodes within complex topology. To further investigate this, we conducted an in-depth analysis of the effects of varying walk numbers (Fig. 2(a)) and walk lengths (Fig. 2(b)). Our evaluation is based on the undirected SRW, detailed in Appendix A.1. We categorized the top 50% homophilous nodes as homophily, while the remaining nodes were classified as heterophily.
Key Insight 2: From the perspective of walk number, higher walk numbers can facilitate the prediction of heterophilous nodes. Despite an increase in walk number providing no significant performance boost for homophilous nodes, it substantially enhances predictions for heterophilous nodes. Specifically, with fewer walks, low-homophily nodes fail to capture sufficiently structural insights. Consequently, increasing the number of walks allows for a more comprehensive exploration of a node’s intricate topological contexts. Conversely, nodes with high homophily exhibit a more uniform neighborhood structure, which can be accurately represented with fewer walks, as their local context is abundant in high-quality information. However, excessive walks in such case raise concerns regarding over-smoothing and time and space complexity.
Key Insight 3: From the perspective of walk length, extending walk length positively affects nodes with high homophily but negatively impacts those with low homophily. The underlying cause is rooted in the varying informational landscape depending on the homophily level of nodes. Nodes with low homophily are often surrounded by a noisy environment, where longer walks risk incorporating excessive irrelevant information and distracting features. Conversely, nodes with high homophily are typically surrounded by a more cohesive and informative neighborhood. In such contexts, longer walks are more beneficial as they include a broader spectrum of relevant information, providing the model with a more comprehensive view of the node’s surroundings.
@@
Inspired by the above key insights, in this paper, we propose Directed Random Walk (DiRW), which can be viewed as a plug-and-play directed walk strategy for existing DiPathGNNs or a new learning paradigm for digraphs. Specifically, DiRW first designs a direction-aware walk strategy to identify potential relationships between the current node and its generalized neighbors through walk sequences (motivated by Key Insight 1). Subsequently, DiRW fine-tunes the walk numbers and walk lengths on a node-adaptive basis, thereby providing a rich information source for representing the current node (motivated by Key Insight 2 & 3). After that, DiRW employs a node-wise learnable path aggregator to efficiently represent node embeddings, which breaks the fixed neighborhood constraints of traditional message-passing mechanisms, allowing the exploration of higher-order node semantic consensus.
@@
Our contributions. (1) Novel Perspective. To the best of our knowledge, this paper is the first to propose the concept of DiPathGNN and summarize its advantages. Furthermore, we emphasize the critical need for fine-grained walking strategies in digraphs, offering valuable empirical studies. (2) Plug-and-play Strategy. We introduce DiRW for digraphs, which customizes walk probabilities, length, and number for each node. It can be seamlessly integrated with any DiPathGNN to enhance performance. (3) New DiGNN. DiRW is also regarded as a new learning architecture for digraphs, consisting of an optimized path sampler in the pre-processing stage and a well-designed node-wise path aggregator in the learning process. (4) SOTA Performance. Experiments demonstrate that DiRW achieves an improvement of as a plug-and-play strategy in node classification and as a novel DiGNN in link prediction.
@@ 2. Preliminaries
2.1. Notation
We consider a digraph with nodes and edges. Each node has a feature vector of size and a one-hot label of size , with the feature and label matrices represented as and , respectively. The digraph can be described by an asymmetrical adjacency matrix , where if , and otherwise. Suppose is the labeled set, and the goal of the semi-supervised node classification task is to predict the labels for nodes in the unlabeled set with the supervision of .
@@ 2.2.1. Spatial-based Methods
To capture the asymmetric topology of digraphs, some spatial-based methods follow the strict symmetric message passing paradigm in the undirected setting (Huang et al., 2021; Frasca et al., 2020; Xu et al., 2018; Zhang et al., 2022). However, it is crucial to account for the directionality of the edges when aggregating messages. Specifically, for the current node , the model learns weights to combine the representations of its out-neighbors and in-neighbors independently:
Building upon this strategy, recent advances in digraph GNNs have further refined the message passing scheme to capture the inherent directionality of the digraph. For instance, DGCN (Tong et al., 2020b) incorporates both first- and second-order neighbor proximity into the message aggregation process. DIMPA (He et al., 2022) expands the receptive field of by aggregating features from -hop neighborhoods at each layer, enabling the capture of long-range dependencies. Inspired by the 1-WL graph isomorphism test, NSTE (Kollias et al., 2022) tailors the message propagation to the directed nature of the graph. DiGCN (Tong et al., 2020a) leverages neighbor proximity to increase the receptive field and proposes a digraph Laplacian based on personalized PageRank. ADPA (Sun et al., 2023) adaptively explores appropriate directed patterns to conduct weight-free graph propagation and employs two hierarchical node-wise attention mechanisms to obtain optimal node embeddings.
@@ 2.2.2. Spectral-based Methods
Spectral-based approaches for the DiGNN depart from the strict symmetric message passing used for undirected graphs (He et al., 2021; Wang and Zhang, 2022). The core of spectral-based DiGNNs is to leverage a symmetric or conjugated digraph Laplacian , which is constructed based on the directed adjacency matrix . This symmetric Laplacian allows the application of spectral convolution operations, which can be formally represented as a function of the eigenvalues and eigenvectors of . Specifically, the layer-wise node embeddings are computed via a first-order approximation of Chebyshev polynomials, leveraging the spectral decomposition of the symmetric digraph Laplacian.
@@ 2.3. Path-based Graph Neural Networks
PathGNNs offer an effective approach to capture the intricate patterns within graphs by sampling and aggregating information along paths. For instance, GeniePath (Liu et al., 2019) introduces an adaptive path layer that intelligently navigates the exploration of both the breadth and depth of the node’s receptive fields. SPAGAN (Yang et al., 2021) delves into the topological information of graphs by leveraging the shortest paths and applying path-based attention mechanisms to obtain node embeddings. PathNet (Sun et al., 2022) utilizes a maximal entropy-based random walk strategy to capture the heterophily information while preserving valuable structural information. RAWGNN (Jin et al., 2022) employs Node2Vec (Grover and Leskovec, 2016) to simulate both BFS and DFS, thereby capturing both homophily and heterophily information. PathMLP (Xie et al., 2023) designs a similarity-based path sampling strategy to capture smooth paths containing high-order homophily. The detailed descriptions of PathGNNs can be found in Appendix. A.8. We have also profoundly revealed that PathGNNs can be considered as a generalized and lightweight graph attention network in Appendix. A.9.
@@ 3. Method
Building upon the key insights discussed in Sec. 1, we now present our DiRW model, which is composed of two principal components: the optimized path sampler and the node-wise learnable path aggregator. The architecture of the model is depicted in Fig. 3.
Specifically, drawing upon Key Insight 1, DiRW is initiated with a direction-aware path sampling strategy in Sec. 3.1.1. This strategic approach determines the destinations for each walk step, thereby significantly augmenting the capacity to encapsulate the intricacies of directed topological structures. Furthermore, we introduce a multi-scale and multi-order walk probability scheme in Sec. 3.1.2, which models high-order and first-order homophily from both topological structure and node profiles. Guided by Key Insight 3, we customize the walk length for each path in Sec. 3.1.3. By introducing the homophily entropy, we evaluate the quality of walk sequences and terminate the walk when sequence quality reaches an optimal threshold. In Sec. 3.1.4, we propose a weight-free aggregation mechanism for the path embedding, which preserves the rich information encoded within the sequential order. Inspired by Key Insight 2, we tailor walk numbers in Sec. 3.1.5. By assessing the informational yield of sampled walks, we implement a stopping criterion that ensures the collection of ample yet efficient information, thereby optimizing the model’s performance. In the model learning phase, we integrate an attention mechanism to learn the importance of each walk sequence in Sec. 3.2.1, which prevents the aggregation of ineffective node information. Culminating in Sec. 3.2.2, we utilize a linear layer to distill the final node representations, which are applied to node classification and link prediction.

### v3

@@ Abstract.
Recently, graph neural network (GNN) has emerged as a powerful representation learning tool for graph-structured data. However, most approaches are tailored for undirected graphs, neglecting the abundant information in the edges of directed graphs (digraphs). In fact, digraphs are widely applied in the real world and confirmed to address heterophily challenges. Despite recent advancements, existing spatial- and spectral-based DiGNNs have limitations due to their complex learning mechanisms and reliance on high-quality topology, resulting in low efficiency and unstable performance. To address these issues, we propose Directed Random Walk (DiRW), a plug-and-play strategy for most spatial-based DiGNNs and also an innovative model which offers a new digraph learning paradigm. Specifically, it utilizes a direction-aware path sampler optimized from the perspectives of walk probability, length, and number in a weight-free manner by considering node profiles and topologies. Building upon this, DiRW incorporates a node-wise learnable path aggregator for generalized node representations. Extensive experiments on 9 datasets demonstrate that DiRW: ➀ enhances most spatial-based methods as a plug-and-play strategy; ➁ achieves SOTA performance as a new digraph learning paradigm. The source code and data are available at https://github.com/dhsiuu/DiRW.
@@ 1. Introduction
Graph neural networks (GNNs) have been widely used across node- (Wu et al., 2019; Wang and Leskovec, 2020; Li et al., 2024b), link- (Zhang and Chen, 2018; Tan et al., 2023; Li et al., 2024a), and graph-level tasks (Liang et al., 2023; Luo et al., 2023; Thost and Chen, 2021) and achieve satisfactory performance. Therefore, this graph-based deep learning technique holds great potential for applications, such as recommendation (Su et al., 2024; Yang et al., 2023), financial analysis (Balmaseda et al., 2023; Hyun et al., 2023; Qiu, 2023), and healthcare (Pfeifer et al., 2022; Ahmed et al., 2022). Despite their effectiveness, existing methods often overlook edge direction in natural graphs, leading to inevitable information loss and limited performance upper bound.
@@
Compared to undirected representations, digraphs are crucial for modeling complex real-world topologies (e.g., web flow monitoring and bioinformatics), capturing more intricate node relationships. Additionally, the recently proposed A2DUG (Maekawa et al., 2023), Dir-GNN (Rossi et al., 2023) and ADPA (Sun et al., 2024) reveal a key insight: Edge direction offers a new perspective for addressing the topological heterophily challenges that plague graph learning. Despite growing attention, DiGNNs are still in their infancy and face the following inherent limitations: ➀ Spatial-based methods often stack multiple convolution layers with separate learnable parameters for out-edges and in-edges, resulting in over-smoothing concerns (Yan et al., 2022) and high computational costs (He et al., 2022; Kollias et al., 2022; Zhou et al., 2022). ➁ Spectral-based methods heavily rely on high-quality directed topology (Zhang et al., 2021a; Tong et al., 2020a), and without this, extreme eigenvalues inevitably lead to sub-optimal performance. Strict theoretical assumptions also limit practical deployment in complex scenarios. Therefore, developing a more efficient paradigm for digraph learning is urgent.
@@
To enhance the usability of DiGNNs, this paper focuses on spatial-based methods and proposes a novel (directed) path-based learning mechanism (Di)PathGNNs. As we all know, the entanglement of homophily and heterophily, where connected nodes exhibit intricate feature distributions and labels, has recently posed a significant challenge (Ma et al., 2021; Luan et al., 2022; Zheng et al., 2022; Platonov et al., 2023). Researchers strive to achieve robust learning within this complex topology. In this context, compared to the traditional message-passing (i.e., neighbor expansion with strict spatial symmetry that disregards edge direction), we highlight the following advantages of DiRW to further clarify the motivation of our study: ❶ Edge Direction and Node Order. The core of DiPathGNNs lies in performing well-designed random walks for each node, treating the paths as node-wise sequences. Advantage: DiRW fully considers edge direction and preserves the order of nodes within the walking paths, which is crucial for capturing structural insights. ❷ Adaptive Expansion of Node Receptive Fields. DiPathGNNs adaptively extend hop-based neighbors to path-based neighbors by considering the characteristics of random walk, incorporating more homophilous signals. Advantage: DiRW ensures message aggregation among nodes with the same label, equivalent to data augmentation and highlighting label-specific positive signals for prediction. ❸ Path-based Message Aggregation. Based on the above homophily-aware paths, DiPathGNNs aggregate these messages for predicting each node through a learnable mechanism. Advantage: DiRW facilitates the dense aggregation of direction-aware homophilous information, thus enhancing node prediction.
@@
Despite recent advancements in PathGNNs (Jin et al., 2022; Xie et al., 2023; Sun et al., 2022), their walk strategies are tailored for undirected graphs, lacking generalizability. Furthermore, the complex relationships inherent in directed topology pose unique challenges to the naive random walks, necessitating further investigations to develop fine-grained walk rules. To further elucidate, we present empirical analyses in Fig. 1-2 that demonstrate these limitations and offer key insights.
@@
◆Limitation 1: Neglect of edge directions. PathGNNs often fail to account for the edge direction. Utilizing directed simple random walks (DiSRW) on digraphs inevitably leads to interruptions when encountering nodes lack outgoing edges. To investigate and visualize this limitation, we employ DiSRW and initiate a walk from each node on four digraphs. As the walk length increases, we record the proportion of complete paths out of the total sequences, as depicted in Fig. 1(a). To mitigate the impact of cycles within the digraph, we also devise an DiSRW that excludes cycles and conduct the same experiment, as shown in Fig. 1(b).
@@
✰Key Insight 1: Walking strictly along the edge directions leads to severe walk interruption problems. The non-strongly connected nature of digraphs causes most walks to show a sharp decline in complete paths after just five steps, indicating that they fail to gather extensive information beyond the immediate neighborhood of the starting node. Additionally, removing the influence of cycles results in an even greater decline in uninterrupted sequences.
@@
◆Limitation 2: Coarse-grained walk strategies. PathGNNs treat walk number and length as hyperparameters, uniformly applying them across all nodes. This one-size-fits-all approach overlooks the distinct contexts of nodes within complex topology. To further investigate this, we conducted an in-depth analysis of the effects of varying walk numbers (Fig. 2(a)) and walk lengths (Fig. 2(b)) on node classification. We utilize undirected SRW as the sampling strategy, and utilize two MLPs to generate path embeddings from the nodes along the sampled paths and aggregate node embeddings from multiple paths associated with the same node. Node homophily is calculated as the ratio of the number of first-order neighboring nodes sharing the same label to the total number of first-order neighbors, with the top 50% of nodes by homophily scores classified as homophilous nodes and the remains as heterophilous nodes.
✰Key Insight 2: Higher walk numbers facilitate heterophilous nodes, while longer walk lengths benefit homophilous nodes. Increasing walk numbers enhances predictions for heterophilous nodes by enabling a comprehensive exploration of their complex topologies. In contrast, homophilous nodes can be accurately represented with fewer walks due to their uniform neighborhood structure. Excessive walks in such case raise concerns regarding over-smoothing and time and space complexity. Regarding walk length, heterophilous nodes are often surrounded by noisy environments, where longer walks risk incorporating irrelevant information. Conversely, homophilous nodes benefit from longer walks as they gather a broader range of relevant information, offering a clearer surrounding views.
@@
Inspired by the above key insights, we propose Directed Random Walk (DiRW), which can be viewed as a plug-and-play directed walk strategy for PathGNNs or a new digraph learning paradigm. Specifically, DiRW first designs a direction-aware walk strategy to identify potential neighbor relationships (motivated by Key Insight 1). It then fine-tunes walk numbers and lengths on a node-adaptive basis (motivated by Key Insight 2). Finally, DiRW employs a node-wise learnable path aggregator to represent node embeddings, breaking the constraints of message-passing mechanisms.
@@
Our contributions. ➀ Novel Perspective. This paper is the first to introduce DiPathGNN and highlight its advantages, emphasizing the need for fine-grained walking strategies in digraphs through valuable empirical studies. ➁ Plug-and-play Strategy. We present DiRW, which customizes walk probabilities, lengths, and numbers for each node, seamlessly integrating with any DiPathGNN to enhance performance. ➂ New DiGNN. DiRW also serves as a new learning architecture for digraphs, featuring a pre-processed optimized path sampler and a well-designed node-wise learnable path aggregator. ➃ SOTA Performance. Experiments demonstrate that DiRW improves node classification by as a plug-and-play strategy and enhances link prediction by as a novel DiGNN.
@@ 2. Preliminaries
2.1. Notation and Problem Formulation
We consider a digraph with nodes and edges, where each node has a feature of size and a one-hot label of size . The feature and label matrices represented as and . The is described by an asymmetrical adjacency matrix , where if , and otherwise. The goal of the semi-supervised node classification is to predict the labels for unlabeled nodes with the supervision of labeled nodes.
@@ 2.2.1. Spatial-based Methods
To capture the asymmetric topology of digraphs, some spatial-based methods follow the strict symmetric message-passing paradigm in the undirected setting (Huang et al., 2021; Frasca et al., 2020; Xu et al., 2018). However, it is crucial to account for the directionality of the edges when aggregating messages. Specifically, for the current node , the model learns weights to combine the representations of its out-neighbors and in-neighbors independently:
where the Propagation function gathers and distributes feature from neighbors, and the Aggregation function combines the feature depending on the learnable parameter matrix . Building upon this, recent advances in DiGNNs further refine the message-passing scheme to capture the inherent directionality of the digraph. DGCN (Tong et al., 2020b) incorporates both first- and second-order neighbor proximity into the message aggregation process. DIMPA (He et al., 2022) expands the receptive field by aggregating features from -hop neighborhoods. Inspired by the 1-WL graph isomorphism test, NSTE (Kollias et al., 2022) tailors the message propagation to the directed nature of the graph. DiGCN (Tong et al., 2020a) leverages neighbor proximity to increase the receptive field and proposes a digraph Laplacian based on personalized PageRank. ADPA (Sun et al., 2024) adaptively explores directed patterns to conduct weight-free propagation and employs two hierarchical node-wise attention mechanisms to learn representations.
@@ 2.2.2. Spectral-based Methods
Spectral-based approaches for the DiGNN depart from the strict symmetric message-passing used for undirected graphs (He et al., 2021; Wang and Zhang, 2022). The core of spectral-based DiGNNs is to leverage a symmetric or conjugated digraph Laplacian , which is constructed based on the directed adjacency matrix . This symmetric Laplacian allows the application of spectral convolution operations, which can be formally represented as a function of the eigenvalues and eigenvectors of . Specifically, the layer-wise node embeddings are computed via a first-order approximation of Chebyshev polynomials, leveraging the spectral decomposition of the symmetric digraph Laplacian.
@@ 2.3. Path-based Graph Neural Networks
PathGNNs offer an effective approach to capture the intricate graph patterns by sampling and aggregating information along paths. For instance, GeniePath (Liu et al., 2019) introduces an adaptive path layer that navigates the exploration of both the breadth and depth of the node’s receptive fields. SPAGAN (Yang et al., 2021) leverages the shortest paths and applies path-based attention mechanisms to obtain node embeddings. PathNet (Sun et al., 2022) utilizes a maximal entropy-based random walk strategy to capture the heterophilous and structural information. RAWGNN (Jin et al., 2022) employs Node2Vec (Grover and Leskovec, 2016) to simulate both BFS and DFS, capturing both homophily and heterophily information. PathMLP (Xie et al., 2023) designs a similarity-based path sampling strategy to capture smooth paths containing high-order homophily.
@@ 3. Method
Building upon the key insights discussed in Sec. 1, we now present our DiRW model, which is composed of two principal components: the optimized path sampler and the node-wise learnable path aggregator. The architecture of the model is depicted in Fig. 3. Specifically, drawing upon Key Insight 1, DiRW is initiated with a direction-aware path sampling strategy in Sec. 3.1.1. Furthermore, we introduce a multi-order walk probability scheme in Sec. 3.1.2. Guided by Key Insight 2, we evaluate the quality of walk sequences by introducing the homophily entropy to customize the walk length in Sec. 3.1.3. In Sec. 3.1.4, we propose a weight-free aggregation mechanism for the path embedding. We also tailor walk numbers in Sec. 3.1.5 by assessing the information richness of sampled walks. In the learning phase, we integrate an attention mechanism to learn the importance of each walk sequence in Sec. 3.2.1. Culminating in Sec. 3.2.2, we utilize a linear layer to distill the final node representations, applied to node classification and link prediction.